OpenStudy (anonymous):

The cost of producing x units in a manufacturing process is given by the function C(x)= 1.25x+65. the revenue obtained from selling x units is given by R(x)= 2.75x-0.0025x^2. determine the profit as a function of the number of units sold if P=R-C.

OpenStudy (anonymous):

hahahahah nadine

OpenStudy (anonymous):

i'm busy facebook helping shauna hehe

OpenStudy (anonymous):

R(x) - C(x) = P(x) 2.75x - 0.0025x^2 - 1.25x + 65 = P(x)